OpenStudy (anonymous):

write an equation in point-slope form of the line through the points j(4 1) with slope -4

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(4\quad ,&1) \end{array} \\\quad \\ slope = m= -4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values}\)